A kitchen oven has an interior volume V of 0.12 cubic meter. The initial pressure P inside the oven equals the outside pressure of 1000 mbaqanga. The air in the oven is then heated from an initial temperature T of 70 degree Fahrenheit to a baking temperature T of 425 degree Fahrenheit.
a) If the oven were perfectly sealed so that no air escaped, what would the pressure be insides the oven when it reached its final temperature?
b) If the door has an area of 0.2 meter squares, what would the net force acting on the door be if someone opened it without relieving the pressure first?
c) If the oven is not sealed (normal case), air will leak out as the air expands from heating so that the inside pressure always equals the outside pressure. Assuming that the air is oneway (no outside air enters the oven), what fraction of the original mass of air in the oven is lost due to expansion?
1
Expert's answer
2019-09-16T04:05:34-0400

a) pV=nRT. n=pV/RT. p= nRT/ V

n= 1000 x 0.12 / 0.082 x 294,261= 5 moles.

p= 5 x 0.082 x 491.483 / 0.12= 1679.23358


b) p=F/S. F= p x S = 1679.23358 x 0.2 = 335.846716 N


c) pV= mRT / M. m= pVM / RT= (1000 x 0.12 x 29) / (0.082 x 294.261) = 144.222389 g.